# Multimodal Financial Foundation Models (MFFMs): Progress, Prospects, and Challenges

### v1

@@ Abstract.
Financial Large Language Models (FinLLMs), such as open FinGPT and proprietary BloombergGPT, have demonstrated great potential in select areas of financial services. Beyond this earlier language-centric approach, Multimodal Financial Foundation Models (MFFMs) can digest interleaved multimodal financial data, including fundamental data, market data, data analytics, macroeconomic, and alternative data (e.g., natural language, audio, images, and video). In this position paper, presented at the MFFM Workshop joined with ACM International Conference on AI in Finance (ICAIF) 2024, we describe the progress, prospects, and challenges of MFFMs. This paper also highlights ongoing research on FinAgents in the SecureFinAI Lab [^1] https://openfin.engineering.columbia.edu/ at Columbia University. We believe that MFFMs will enable a deeper understanding of the underlying complexity associated with numerous financial tasks and data, streamlining the operation of financial services and investment processes.
Github Repo: https://github.com/Open-Finance-Lab/Awesome-MFFMs/
copyright: acmlicensedjournalyear: 2024copyright: acmlicensedconference: International Workshop on Multimodal Financial Foundation Models (MFFMs) at 5th ACM International Conference on AI in Finance; Nov. 15; NY, USAbooktitle: International Workshop on Multimodal Financial Foundation Models (MFFMs) at 5th ACM International Conference on AI in Finance (MFFM at ICAIF ’24), November 14–17, 2024, Brooklyn, NY, USA
Figure 1. An overview. Multimodal FinData (left block) is ubiquitous in FinApps (middle block), such as search, tutor, robo-advisor, compliance, auditing, trading, etc. However, several major challenges (right block) call for immediate actions, in order to achieve FinAI readiness and governance.
@@ 1. Introduction
Large language models (LLMs) have demonstrated remarkable proficiency in understanding and generating human-like texts. FinLLMs, such as open FinGPT (Liu et al., 2023, 2024a; Yang et al., 2023; Zhang et al., 2023a; Tian et al., 2024) and proprietary BloombergGPT (Wu et al., 2023), have shown great potential in select areas of financial services (Kong et al., 2024a, b) . Beyond this earlier language-centric approach, Multimodal Financial Foundation Models (MFFMs) can process interleaved multimodal financial data, including fundamental data, market data, data analytics, macroeconomics, and alternative data (e.g. natural language, audio, image and video). Multimodal financial data has unique characteristics, such as dynamic, both structured and unstructured forms, and comes in varying formats (e.g., charts, graphs, Web APIs, Excel spreadsheets, SEC filings, XBRL filings (Han et al., 2024), and SQL data (Zhao et al., 2024)).
@@
On the journey toward widespread adoption of MFFMs, several challenges remain, including increasing concerns related to reproducibility, transparency, privacy, and ethics. First, many existing LLMs function as black boxes, posing challenges in comprehending their operations and ensuring fairness. Another two major challenges are “model cannibalism” and “openwashing.” Many models are largely trained and released without transparency in mind, e.g., Claude 3.5 Sonnet. Many supposedly ”novel” models may exploit labels from existing LLMs (e.g., GPT-4o) and perform supervised learning, referred to as “model cannibalism.” As a result, MFFMs are opaque in decision-making. They give rise to various challenges, such as inadequate transparency concerning the training data, deficiencies in combating models’ inherent biases, safety and security issues, and adversarial attacks (e.g., backdoor attacks).
@@
Recently, there have been many “openwashing” behaviors, where the LLM weights are marketed as ”open” but under restricted licenses, while OSI-approved licenses (e.g., Apache License 2.0 and MIT License) are preferable. A research alliance among Columbia University, Oxford University, and the Linux Foundation proposed the Model Openness Framework (White et al., 2024) that classifies model openness by ranking models according to $1616$ components. This framework offers guidance to researchers and model producers for promoting transparency and reproducibility. MFFMs that comply with this framework would promote reproducibility and adoption across the finance industry, such as Open FinLLMs (Xie et al., 2024b). For financial institutions, it provides clear guidelines for new models to become commercially suitable without restrictions.
@@
In this position paper, presented at the MFFM Workshop [^2] MFFM Website: https://sites.google.com/view/iwmffm2024/home?authuser=1 jointly held with ACM International Conference on AI in Finance (ICAIF) 2024, we describe the progress, prospects, and challenges of MFFMs. This paper also highlights ongoing research on FinAgents in the SecureFinAI Lab at Columbia University. We first list multimodal financial data and data-centric approach in the financial domain (left block of Fig. 1). Then, we describe multimodal financial applications (middle block of Fig. 1). We envision that the AI agent is a promising solution to build multiple financial applications. Using the Model Openness Framework, licenses (e.g. OpenMDW [^3] https://openmdw.ai/) and agent protocols (e.g. Model Context Protocol, Agent2Agent Protocol) can build financial Agentic AI ecosystems. However, several major challenges (right block of Fig. 1) call for immediate actions in order to achieve financial AI readiness. The associated challenges are proprietary data constraints, training and inference costs, regulatory complexities, reasoning capacity, and the need for robust benchmarks and a guardrail framework to address misinformation and data biases. We believe that MFFMs will enable promising financial tasks and data analysis, streamlining the operation of financial services.
@@
Related Work: There are several related surveys on FinLLMs, as outlined in Table 1. Li et al. (2023) first reviewed the approach employing LLMs in finance. Ding et al. (2024) summarized the performance of LLM-based agents in financial trading tasks. Lee et al. (2024) reviewed FinLLMs from a benchmark perspective. Kong et al. (2024a) and Kong et al. (2024b) further summarize recent advancements in FinLLMs and discuss their various application scenarios. Despite these efforts, these surveys lack a comprehensive review of multimodal financial data, applications, and models. Furthermore, there is still a need for an in-depth analysis concerning the opportunities and challenges of applying LLMs in finance, including their current readiness level.
@@
  • To the best of our knowledge, this is the first comprehensive survey of multimodal financial foundation models (MFFMs). We summarize three aspects, multimodal financial data, applications, and model exploration.
  • For multimodal financial data, we emphasize a data-centric approach. For applications, we point out the trend of agentic AI ecosystem, which is enabled by open models and the model context protocol.
  • We compare and contrast MFFMs with LLMs, FinLLMs, and MM-LLMs. Our aim is to offer readers a holistic view of MFFM development and help readers understand the current progress and future prospects.
@@
  • We discuss the ethical challenges of MFFMs’s readiness, including the hallucination and misinformation that may arise from real-world use scenarios, as well as ethical issues.

### v2

@@ Abstract.
Financial Large Language Models (FinLLMs), such as open FinGPT and proprietary BloombergGPT, have demonstrated great potential in select areas of financial services. Beyond this earlier language-centric approach, Multimodal Financial Foundation Models (MFFMs) can digest interleaved multimodal financial data, including fundamental data, market data, data analytics, macroeconomic, and alternative data (e.g., natural language, audio, images, and video). In this position paper, presented at the MFFM Workshop joined with the ACM International Conference on AI in Finance (ICAIF) 2024, we describe the progress, prospects, and challenges of MFFMs. This paper also highlights ongoing research in the SecureFinAI Lab [^1] https://openfin.engineering.columbia.edu/ at Columbia University and summarizes the FinLLM Exploration meetings at FinOS of Linux Foundation. MFFMs will enable users to better understand the underlying complexity associated with numerous financial tasks and data, simplifying the operation of financial services and investment processes.
Github: https://github.com/Open-Finance-Lab/Awesome-MFFMs/
copyright: acmlicensedjournalyear: 2024copyright: acmlicensedconference: International Workshop on Multimodal Financial Foundation Models (MFFMs) at 5th ACM International Conference on AI in Finance; Nov. 15; NY, USAbooktitle: International Workshop on Multimodal Financial Foundation Models (MFFMs) at ACM International Conference on AI in Finance (MFFM at ICAIF), Nov. 14–17, 2024,
Figure 1. An overview. Multimodal financial data (left block) is ubiquitous in real life, while financial applications (middle block) require financial agents to possess multimodal capabilities, such as search, tutor, robo-advisor, compliance, auditing, trading, and low-code development. However, the urgent need for MFFM models (right block) call for immediate actions, in order to achieve FinAI readiness and governance.
@@ 1. Introduction
Generative AI is witnessing the rise of foundation models (e.g., transformer neural network, diffusion models) trained on massive data that can be adapted to a wide range of downstream tasks (Bommasani et al., 2021). Recently, large language models (LLMs) have demonstrated remarkable proficiency in understanding and generating human-like texts. FinLLMs, such as open FinGPT (Liu et al., 2023b, 2024a; Yang et al., 2023; Zhang et al., 2023a, b; Tian et al., 2024) and proprietary BloombergGPT (Wu et al., 2023), have shown great potential in select areas of financial services (Kong et al., 2024a, b). Beyond this earlier language-centric approach, Multimodal Financial Foundation Models (MFFMs) can process interleaved multimodal financial data, including fundamental data, market data, data analytics, macroeconomics, and alternative data (e.g. natural language, audio, visual). Multimodal financial data has unique characteristics, such as dynamic, both structured and unstructured forms, and comes in varying formats (e.g., charts, graphs, Web APIs, Excel spreadsheets, SEC filings (Islam et al., 2023), XBRL filings (Han et al., 2024), and SQL data (Zhao et al., 2024)).
@@
On the journey toward widespread adoption of MFFMs, several challenges remain, including increasing concerns related to reproducibility, transparency, privacy, and ethics. First, many existing LLMs function as black boxes, posing challenges in comprehending their operations and ensuring fairness. Another two major challenges are “model cannibalism” and “openwashing.” Many models are largely trained and released without transparency in mind, e.g., the Claude 3.5 Sonnet. Many supposedly ”novel” models may exploit labels from existing LLMs (e.g., GPT-4o) and perform supervised learning, referred to as “model cannibalism.” As a result, MFFMs are opaque in decision-making. They give rise to various challenges, such as inadequate transparency in training data, deficiencies in combating models’ inherent biases, safety and security issues, and adversarial attacks (for example, backdoor attacks).
@@
Recently, there have been many “openwashing” behaviors, where the LLM weights are marketed as ”open” but under restricted licenses, while OSI-approved licenses (e.g., Apache License 2.0 and MIT License) are preferable. A research alliance between Columbia University, Oxford University, and the Linux Foundation proposed the Model Openness Framework (White et al., 2024) that classifies model openness by ranking models according to $1616$ components. This framework offers guidance to researchers and model producers for promoting transparency and reproducibility. MFFMs that comply with this framework would promote reproducibility and adoption across the finance industry, such as Open FinLLMs (Xie et al., 2024b). For financial institutions, it provides clear guidelines for new models to become commercially suitable without restrictions.
@@
In this position paper, presented at the MFFM Workshop [^2] MFFM Workshop: https://sites.google.com/view/iwmffm2024/home?authuser=1 jointly held with ACM International Conference on AI in Finance (ICAIF) 2024, we describe the progress, prospects, and challenges of MFFMs. This paper also highlights ongoing research in the SecureFinAI Lab at Columbia University and and summarizes the FinLLM Exploration meetings at FinOS of Linux Foundation. We first list multimodal financial data and data-centric approach in the financial domain (left block of Fig. 1). Then, we describe multimodal financial applications (middle block of Fig. 1). We envision that FinAgent is a promising solution to build multiple financial applications. Following the Model Openness Framework, open licenses (e.g. OpenMDW [^3] https://openmdw.ai/) and agent protocols (e.g. Model Context Protocol, Agent2Agent Protocol) enables the blooming of an agentic FinAI ecosystem. However, several major challenges (right block of Fig. 1) call for immediate actions in order to achieve FinAI readiness. The associated challenges are proprietary data constraints, training and inference costs, regulatory complexities, reasoning capacity, and the need for robust benchmarks and a guardrail framework to address misinformation and data biases. We believe that MFFMs will enable promising financial tasks and data analysis, streamlining the operation of financial services.
@@
Types Text Audio Image Video Numbers Tabular Chart Time-Series
Earnings Conference Calls (ECC)
Monetary Policy Calls (MPC)
Climate Data
Financial News
Market Data
Financial Reports
Financial curriculum and certifications
Table 2. Overview of multimodal financial data.
Related Work: Several related surveys are given in Table 1. Li et al. (2023) first reviewed the approach employing LLMs in finance. Ding et al. (2024) summarized the performance of LLM-based agents in financial trading tasks. Lee et al. (2024) reviewed FinLLMs from a benchmark perspective. Kong et al. (2024a) and Kong et al. (2024b) further summarize recent advancements in FinLLMs and discuss their various application scenarios. Despite these efforts, the rapidly evolving nature of the field necessitates an updated and thorough review of multimodal financial data, applications, and models. Furthermore, there is still a need for an in-depth analysis concerning the opportunities and challenges of applying LLMs in finance, including their current readiness level.
@@
  • To the best of our knowledge, this is the first comprehensive survey of MFFMs. We summarize three aspects, multimodal financial data, applications, and model exploration.
  • For multimodal financial data, we emphasize a data-centric approach. For applications, we point out the promising era of an agentic AI ecosystem with various types of FinAgents, which is enabled by open models and agent protocols.
  • We compare and contrast MFFMs with LLMs, FinLLMs, and MM-LLMs. Our aim is to offer readers a holistic view of the MFFM development lifecycle and help readers understand current progress and future prospects.
@@
  • We discuss ethical challenges of MFFMs’s readiness, including the hallucination and misinformation in use scenarios, as well as the importance of building a guardrail framework.